The sum of the squares of two consecutive positive numbers is 265. Find both numbers.
Answer by jorel1380(3719)

x2+(x+1)2=265
x2+x2+2x+1=265
2x2+2x-264=0
x2+x-132=0
(x+12)(x-11)=0
x=-12 or ll
Throwing away the negative answer, we get the two consecutive integers to be 11 and 12..